OpenStudy (anonymous):

Angle DEF and angle FEG are supplementary. The measure of angle DEF = (9x +1) and the measure of angle FEG = (8x + 9). What is the measure of angle FEG?

OpenStudy (anonymous):

angle DEF + angle EFG = 180 SOLVE FOR X

OpenStudy (anonymous):

i got 8 for x

OpenStudy (anonymous):

WRONG!

OpenStudy (anonymous):

this is frustrating i have the problem set up as 9x+1=8x+9 subtracted 8x from both sides

OpenStudy (anonymous):

9x+1+8x+9=180

OpenStudy (anonymous):

find x

OpenStudy (anonymous):

can you do it?

OpenStudy (anonymous):

im trying but would i start from the right when subtracting?

OpenStudy (anonymous):

you should have 17x in the left and 170 to the right

OpenStudy (anonymous):

so you would get 1

OpenStudy (anonymous):

x=170/17 = 10

OpenStudy (anonymous):

my answers that are on my hw are 91,89,9,1 degrees

OpenStudy (anonymous):

now you have x=10 angle FEG= 8x+9

OpenStudy (anonymous):

i got 89 as my answer.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i get that now, but i need more he;[ on a few more problems you wouldn't mind helping

OpenStudy (anonymous):

OKEY

OpenStudy (anonymous):

my next question is If the measure of angle T is (123-5x)º, what is the measure of its complement?

OpenStudy (anonymous):

angle T + its complement = 90

OpenStudy (anonymous):

so 123-5x plus 90 minus x

OpenStudy (anonymous):

= 90 - (123-5x) can you simplify that?

OpenStudy (anonymous):

no not actually i though it was suppose to be 90-x

OpenStudy (anonymous):

90-123+5x =(-33+5x) degree

OpenStudy (anonymous):

i got that

OpenStudy (anonymous):

good

OpenStudy (anonymous):

next one

OpenStudy (anonymous):

Suppose JK bisects angle LJM. The measure of angle LJK = (-10x + 3) and the measure of KJM = (-x + 21)º. Find the measure of angle LJM.